OpenStudy (anonymous):

If the diameter of a circle has endpoints P(–10, –2) and Q(4, 6), where is the center of the circle, how long is its radius, and what is the equation for the circle?

OpenStudy (anonymous):

The center is the midpoint of a diameter so use the midpoint formula. C = ([-10 + 4]/2, [-8 + 4]/2) C = (-3, -2) Then use distance formula to get the radius (use either diameter endpoint) r = √( [4 - -3]^2 + [4 - -2]^2) = √(7^2 + 6^2) = √(49 + 36) = √85 The equation of a circle with center (h,k) and radius r is (x - h)^2 + (y - k)^2 = r^2 so (x + 3)^2 + (y + 2)^2 = 85

OpenStudy (lgbasallote):

the radius would be half the distance of the diameter...you do know how to do the distance formula right? do the distance formula..that's your diameter..divide it by 2 that's your radius.. as for the center..it's the midpoint of your diameter

OpenStudy (campbell_st):

seems to be a problem with the centre, in particular the y value

OpenStudy (lgbasallote):

the equation of the circle is just combine all the gathered information from the coordinates of the center and the length of the radius...do you need further hints? @Flapdragon ?

OpenStudy (anonymous):

Thank you! c: It's fine.

OpenStudy (lgbasallote):

glad to help <tips hat>

OpenStudy (anonymous):

myplesure og helping u

OpenStudy (campbell_st):

I think the centre is x (-10 + 4)/2 = -3 y = (-2 + 6)/2 = 2 centre is (-3, 2) radius = \[\sqrt{(-10 +3)^2 (-2 -2)^2} = \sqrt{49 + 16} = \sqrt{65}\] so the equation of the circle is (x +3)^2 + (y -2)^2 = 65